Detector Systems Update MICO March 3, 2008. MICE Detector Systems CKOV u CKOV humidity (HIH4000) and temperature (LM35) sensors in hand for installation.
Published byModified over 5 years ago
Presentation on theme: "Detector Systems Update MICO March 3, 2008. MICE Detector Systems CKOV u CKOV humidity (HIH4000) and temperature (LM35) sensors in hand for installation."— Presentation transcript:
MICE Detector Systems CKOV u CKOV humidity (HIH4000) and temperature (LM35) sensors in hand for installation. u A long range plan for using Lecroy HV4035 HV modules w Camac controller for CKOV in progress. s CAEN supplies used for startup. TOF0 u Assembly is progressing with cosmics testing on single counters u Extra calibration barrette for cosmics trigger in-situ prepared u About 8 defective PMTs have been returned to Hamamatsu for repair s There are discussions for a more aggressive delivery schedule for new PMTs u Tests on laser injection for single barrettes has given good results, s Probably a simple laser diode may be used instead of more complicate and expensive systems u Tests on HV/signal cables (45+45) under way TOF1 trolley u Construction at MIB mechanics workshop, nearly finished
TOF1 u Waiting mainly for new/repaired PMTs delivery TOF2 u Monitoring delivery schedule for PMTs, Bicron Scintillator u Open question: delivery of pure iron (ARMCO) shielding material Tracker u Cryo-systems #1 and #2 now functioning to spec s Assembling problem with Cryo #3 and #4 was discovered –Most likely resulted in poor cool-down performance for these two systems –#3 and #4 will be retrofitted u Readout tests of AFEIIt boards continue. s The Lead EE working on the AFEIIt readout has left Fermilab as of today –Will be replaced with other EEs in the PPD/EED department –However, the group is down one EE and the work load has not changed –This will cause some delay s Estimated shipping to RAL first week in April –Full readout functionality will not be ready at this time, but readout for CR test will be the focus. Left to do: Clock sync to ISIS clock 600 muon/spill readout rate. Will start with 100
Calorimeter u KL s The KL's iron shield frame has been reassembled and the detector has been re-cabled. s Next week will be devoted to the construction of the new trigger counters with their own shields for magnetic field. u SW s No Update Beam Monitors u Large Beam Monitor work is taking longer than expected s Target ship of LBM by 3/13 problematic s Working to remedy –Manpower added u Fermilab Team arrives at RAL on 3/24 for installation and commissioning